OpenStudy (anonymous):

Which numbers are greater than 7 x 10^−4? Choose all answers that are correct. A. 8 x 10^−4 B. 9 x 10^−5 C. 4 x 10^−4 D. 3 x 10^−2 E. 7 x 10^−5

OpenStudy (compassionate):

\[7 \times 10 - 4 = 70 - 4 = 66\] So, in other words, what numbers are greater than 66

OpenStudy (anonymous):

^^^

OpenStudy (anonymous):

A and B

OpenStudy (anonymous):

@randypolle No direct answers.

OpenStudy (compassionate):

@randypolle , please do not give out answers. It hinders the learning experience.

OpenStudy (anonymous):

thank you know i understand

OpenStudy (anonymous):

wait its 7 x 10^-4!!!